An art store sells brushes for $1.80 each and tubes of paint for $2.20 each. Sandra spent $1.20 more on brushes than on paint. She spent a total of $27.60. how many brushes and tubes of paint did she buy?



answer:sandra bought 8 brushes and 6 tubes of paint
need a full break down of how my book came up with these answers

Answer by Mathtut(3670)

let b and p be the number of brushes and tubes of paint respectively
:
1.80b=2.2p+1.2........eq 1
1.80b+2.2p=27.60.....eq 2
:
rewrite eq 1 to b=((2.2p+1.2)/1.8) and plug b's value into eq 2
:
1.8((2.2P+1.2)/1.8)+2.2p=27.60
:
2.2P+1.2+2.2p=27.60
:
4.4p=26.4
:
p=6 tubes of paint
:
b=2.2(6)+1.2/1.8=8 brushes